# Learning Contrastive Feature Representations for Facial Action Unit Detection

### v4

@@ Abstract
Facial action unit (AU) detection has long encountered the challenge of detecting subtle feature differences when AUs activate. Existing methods often rely on encoding pixel-level information of AUs, which not only encodes additional redundant information but also leads to increased model complexity and limited generalizability. Additionally, the accuracy of AU detection is negatively impacted by the class imbalance issue of each AU type, and the presence of noisy and false AU labels. In this paper, we introduce a novel contrastive learning framework aimed for AU detection that incorporates both self-supervised and supervised signals, thereby enhancing the learning of discriminative features for accurate AU detection. To tackle the class imbalance issue, we employ a negative sample re-weighting strategy that adjusts the step size of updating parameters for minority and majority class samples. Moreover, to address the challenges posed by noisy and false AU labels, we employ a sampling technique that encompasses three distinct types of positive sample pairs. This enables us to inject self-supervised signals into the supervised signal, effectively mitigating the adverse effects of noisy labels. Our experimental assessments, conducted on four widely-utilized benchmark datasets (BP4D, DISFA, GFT and Aff-Wild2), underscore the superior performance of our approach compared to state-of-the-art methods of AU detection. Our code is available at https://github.com/Ziqiao-Shang/AUNCE.
Index Terms:
@@
Facial action unit detection, Contrastive learning, Positive sample sampling, Importance re-weighting strategy.I Introduction
@@
Facial Action Coding System (FACS) [1] is a widely employed approach for facial expression coding, which introduced a set of 44 facial action units (AUs) that establish a link between facial muscle movements and facial expressions [2]. AU detection is a multi-label binary classification problem. Each AU represents a kind of label, and activated and inactivated AUs are respectively labeled 1 and 0. Nowadays, AU detection [3, 4, 5, 6, 7, 8, 9] has emerged as a computer vision technique to automatically identify AUs in videos or images. This advancement holds promise for diverse applications, including human-computer interaction [10], emotion analysis [11, 12], and car driving monitoring [13].
@@
In recent years, there has been a prevailing trend in supervised frameworks for AU detection [14, 15, 16, 17, 18, 19, 20]. These studies have predominantly utilized the binary cross-entropy loss to strictly align predicted and ground truth labels, which is often unnecessary for classification tasks. This overemphasis on label alignment imposes excessive demands on the fitting capacity of models, consequently increasing the risk of overfitting. Moreover, these methods commonly focus on encoding entire pixel-level information of AUs, disregarding the fact that capturing subtle feature differences when AUs activate is sufficient. As a result, these learning methods place significant demands on the model complexity to encode excessive redundant information, ultimately limiting the generalization ability of the model.
@@
Representation learning approaches rooted in contrastive learning have demonstrated substantial promise [21, 22, 23, 24, 25] for classification. This approach aims to maximize the consistency of positive samples that are semantically similar while minimizing the consistency of negative samples that are semantically dissimilar [26]. By optimizing the contrastive loss, the mutual information between variables is either explicitly or implicitly optimized [27]. However, the direct integration of contrastive learning into AU detection encounters two principal challenges. The first challenge is the class imbalance issue of each AU type [14]. Most AU classes manifest a substantial imbalance between majority and minority class samples, with inactivated AUs vastly outnumbering their activated counterparts. This imbalance issue influences the classification decision threshold, potentially resulting in the direct classification of the majority of AUs as the negative class. The second challenge pertains to the noisy and false label problem [28]. Due to the subtle feature differences when AUs activate leading to the inherent difficulty in expertly labeling AUs, existing AU datasets may exhibit a few noisy and false labels, significantly undermining the overall fitting ability of models.
@@
To address these challenges, we propose a contrastive learning framework designed to encode discriminative features for AU detection, as depicted in Fig. 1. This approach serves as a replacement for traditional pixel-level learning methods, which also offers a pathway to the lightweight model development within the field of AU detection. The resultant loss is termed AUNCE, resembling the commonly utilized InfoNCE. Moreover, to mitigate the class imbalance issue of each AU type, we apply a negative sample re-weighting strategy, making modifications to the gradient magnitude of minority and majority class samples during back-propagation by adding importance weight to the denominator of the resultant loss. Additionally, in response to the issue of noisy and false labeling in AUs, we introduce a positive sample samplying strategy to augment the supervision signal through self-supervised signals. This is achieved by establishing sampling ratios for various types of positive sample pairs, ensuring that more robust positive samples are derived from data augmentation with the self-supervised signal. The primary contributions of our work are succinctly summarized as follows:
@@
We propose a contrastive learning framework to learn discriminative feature representations for AU detection, and the resultant loss is named AUNCE.
@@
II Related Work
We provide a comprehensive overview of two areas closely connected to our research: deep contrastive learning and contrastive learning for AU detection.
II-A Deep contrastive learning
@@
In recent years, discriminative contrastive learning, a subset of deep self-supervised learning methods [30], has witnessed substantial progress. This approach adheres to the learn-to-compare paradigm [31], wherein a distinction is made between observed data and noisy or erroneous data, preventing the model from engaging in the reconstruction of pixel-level information. Prominent frameworks in this domain include SimCLR by Chen et al. [26], which introduces a base encoder network and a projection head trained to maximize agreement utilizing a contrastive loss. Additionally, MoCo, proposed by He et al. [32], employs a momentum network to maintain a queue of negative samples for efficient contrastive learning. Seyfi et al. [33] and Zhao et al. [34] have both proposed improved contrastive learning frameworks based on MoCo. Chen and He [35] mitigate undesired collapsing solutions, eschewing the use of negative sample pairs through a stop-gradient operation. Despite variations and errors in representation encoders (denoted as ) and similarity measures across different tasks [36, 37], the underlying concept remains consistent: training the representation encoder by optimizing a contrastive loss [38], achieved through contrasting similar pairs () and dissimilar pairs ().
@@
Apart from that, deep contrastive learning not only facilitates the automated extraction of refined feature representations through the comparison of positive and negative sample pairs but also incorporates diverse contrastive losses for optimization. For example, Oord et al. [27] optimized feature extraction networks by maximizing consistency between predicted and actual outcomes of sequence data, introducing the widely-utilized InfoNCE loss, which implicitly or explicitly serves as a lower bound for mutual information. Theoretical analyses concerning the generalization bounds of contrastive learning for classification tasks have been provided by Arora et al. [39]. Khosla et al. [40] extended the contrastive learning concept to supervised settings with a supervised contrastive learning loss, aiming to enhance feature expression using labeled data. Additionally, Chen et al. [41] proposed a semi-supervised contrastive learning algorithm, initially pretraining all data through contrastive learning and subsequently transferring knowledge to new models via distillation learning with the aid of labeled data.
@@
There is a discernible shift in the construction of sample pairs for contrastive learning. In cases where data is labeled or partially labeled, positive samples can be constructed using images belonging to the same category, rather than being confined to identical images. This approach augments the diversity of positive sample pairs, consequently enhancing feature expression. Furthermore, contrastive learning serves as a precursor to supervised learning, enabling model pretraining to significantly enhance the performance of subsequent supervised learning models.
II-B Contrastive learning for AU detection
@@
Contrastive learning was firstly designed as a pretext task [32] for unsupervised or self-supervised learning and was introduced to the domain of AU detection with limited annotations [42, 43, 44, 45, 46]. For example, Niu et al. [42] introduced a semi-supervised co-training approach that utilized contrastive loss to enforce conditional independence between facial representations generated by two Deep Neural Networks (DNNs). Liu et al. [45] introduced a semi-supervised contrastive learning strategy that concurrently combined region learning, contrastive learning, and pseudo-labeling. Additionally, Chang et al. [46] present a knowledge-driven representation learning framework aimed for AU detection, incorporating a contrastive learning module that leverages dissimilarities among facial regions to effectively train the feature encoder.
@@
Subsequent research endeavors extended contrastive learning to supervised settings, aiming at enhancing feature representation for AU detection through the utilization of labeled data [47, 48, 49]. for example, Wu et al. [47] introduced a contrastive feature learning method, enabling a Convolutional Neural Network (CNN) to acquire distinctive features that represent differences between neutral and AU-activated facial images. Chen et al. [49] introduced a supervised hierarchical contrastive learning method (SupHCL) aimed at enhancing AU detection by fostering knowledge consistency across distinct facial images and various AUs. In a similar vein, Suresh et al. [48] proposed a feature-based positive matching contrastive loss, which facilitates learning distances between positives based on the similarity of corresponding AU embeddings.
Different from the above static image-based approaches, some methods predominantly seek to generalize similarity learning between video frames within the discriminative paradigm [50, 51, 52, 53, 54, 55]. One of the most representative methods is Li et al. [28] proposed, which have capitalized on these attributes by formulating a contrastive self-supervised learning method to encode discriminative AU representations from extensive amounts of unlabeled facial videos.
In summary, contrastive learning has demonstrated promising results in various tasks. However, its direct application to the task of detecting subtle feature differences when AUs activate presents challenges: noisy and erroneous annotations, class imbalance in AUs, and difficulty in distinguishing hard negative samples.
III Method
@@
In Section III, we present a contrastive learning framework specifically designed for the task of AU detection, aiming to incentivize the encoder to learn differential information between AUs from each sample pair, rather than focusing on pixel-level information. Subsequently, we confront two additional challenges: 1. Tackling the issue of class imbalance pertaining to each AU type in order to enhance the accuracy of our framework. 2. Addressing the problem of noisy and false labels, aiming to bolster the robustness of our models against erroneously annotated data.
III-A Contrastive learning framework
@@
For the task of AU detection, which falls under the category of multi-label binary classification, the prevalent approach for training models has been the utilization of cross-entropy loss. This loss function measures the disparity between predicted labels and ground truth labels. However, there exists variability in the occurrence rates of different AUs, with some AUs exhibiting low occurrence rates while others display high occurrence rates. Consequently, in contrast to the conventional cross-entropy loss, we employ a weighted cross-entropy loss [14] denoted as
@@
However, the classification methods based on cross-entropy loss suffer from a common issue of being overly constrained, which can lead to overfitting and limited generalization. This constraint becomes particularly problematic in the context of AU detection, where different AUs share a common encoder. For each AU of every sample, the predicted output must satisfy the aforementioned condition, placing a significant burden on the fitting capability of the encoder. To address this challenge, we introduce a contrastive learning framework for AU detection. As shown in Fig. 1. It comprises a positive sample mining module that integrates self-supervised signals with supervised signals, aiming to address the issue of noisy labels resulting from subtle differences in AUs. Additionally, a negative sample weighting module is included to tackle class imbalance and hard negative sample mining challenges. These modules will be discussed in detail in Sections III-B and III-C, respectively. The resultant loss is formulated as:
@@
The contrastive loss distinguishes itself from weighted cross-entropy loss in the following aspects: 1. The weighted cross-entropy loss is computed based on the sample and its corresponding label , requiring the encoding of pixel-level information in order to align with the label. In contrast, the contrastive loss does not involve explicit labels but is computed based on the features of samples from different classes. This incentivizes the encoder to capture the differential information among different class samples, without the need to encode pixel-level information, enabling a lightweight encoder. 2. The weighted cross-entropy loss imposes constraints on the absolute values of the predicted probabilities. However, for classification tasks, this is often unnecessary and can lead to overfitting as well as demanding higher fitting capability from the encoder. The contrastive loss focuses on constraining the similarity difference between positive and negative sample pairs without explicitly specifying their exact values. This is because the minimal value of is 0, which is achieved when for all negative sample index .
To ensure a fair comparison with conventional pixel-level learning paradigms, we utilize the network structure from the first training stage of MEGraph [56] as our contrastive learning encoder, as illustrated in Fig. 2. The input to the encoder consists of an contrastive larning image pair consisting of an anchor and a random sample, and the encoder is optimized by AUNCE only. Our backbone adopts Swin-transformer [57], and the AUs Relationship-aware Node Feature Learning (ANFL) module is composed of two integral components: the AU-specific Feature Generator (AFG) and the Facial Graph Generator (FGG). The AFG independently generates a unique representation for each AU. Building on these representations, the FGG constructs an optimal graph for each facial display to accurately recognize all target AUs. To accomplish this, the FGG ensures that the AFG encodes task-specific associations among AUs into their respective representations.
III-B Negative sample re-weighting strategy
@@
Above equation follows the Boltzmann distribution. As the value of increases, the entropy of the distribution decreases significantly [59], leading to a sharper distribution in the region of high similarity. This has two consequences: Firstly, it imposes larger penalties on hard negative samples that are in close proximity to the anchor data point, enabling the model to learn more precise decision boundaries from subtle feature differences when AUs activate. Secondly, it controls the relative contribution of negative samples to the parameter updates, providing a means to address class imbalance issues.
@@
Specifically, when anchor data point belongs to the majority class, its positive sample belongs to the majority class, while its negative sample belongs to the minority class. In this case, we aim to have smaller gradient magnitudes for the positive samples (majority class) to update fewer parameters, while the negative samples (minority class) should have larger gradient magnitudes to update more parameters. To achieve this, we begin to set , which increases the impact of parameter updates caused by negative samples (minority class). On the other hand, when belongs to minority class, we begin to set to reduce the influence of parameter updates caused by majority class samples. By adjusting the value of , we can control the relative importance of positive and negative samples during the parameter updating, ensuring that the model pays more attention to minority class samples when necessary and balances the gradient magnitudes accordingly. This helps to balance the representation of different AUs in training set and improves the model ability to precisely classify all AU samples.
III-C Positive sample samplying strategy

### v5

@@ Abstract
Existing methods for facial action unit (AU) detection often rely on encoding pixel-level information of the entire face associated with AU, which not only encodes additional redundant information but also leads to increased model complexity and limited generalizability. Additionally, the accuracy of AU detection is negatively impacted by the class imbalance issue of each AU type, and the presence of noisy and false AU labels. In this paper, we introduce a novel contrastive learning framework aimed for AU detection that incorporates both self-supervised and supervised signals, thereby enhancing the learning of discriminative features for accurate AU detection. To tackle the class imbalance issue, we employ a negative sample re-weighting strategy that adjusts the step size of updating parameters for minority and majority class samples. Moreover, to address the challenges posed by noisy and false AU labels, we employ a sampling technique that encompasses three distinct types of positive sample pairs. This enables us to inject self-supervised signals into the supervised signal, effectively mitigating the adverse effects of noisy labels. Our experimental assessments, conducted on four widely-utilized benchmark datasets (BP4D, DISFA, GFT and Aff-Wild2), underscore the superior performance of our approach compared to state-of-the-art methods of AU detection. Our code is available at https://github.com/Ziqiao-Shang/AUNCE.
keywords:
@@
Facial action unit detection , Contrastive learning , Positive sample sampling , Importance re-weighting strategy1 Introduction
@@
Facial Action Coding System (FACS) [1] is a widely employed approach for facial expression coding, which introduced a set of 44 facial action units (AUs) that establish a link between facial muscle movements and facial expressions [2]. AU detection is a multi-label binary classification problem. Each AU represents a kind of label, and activated and inactivated AUs are respectively labeled 1 and 0. Nowadays, AU detection [3, 4] has emerged as a computer vision technique to automatically identify AUs in videos or images. This advancement holds promise for diverse applications, including human-computer interaction, emotion analysis, car driving monitoring, and stuff like that.
@@
In recent years, there has been a prevailing trend in supervised frameworks for AU detection [5, 6, 7, 8, 9]. These studies have predominantly utilized the binary cross-entropy loss to strictly align predicted and ground truth labels, which is often unnecessary for classification tasks. This overemphasis on label alignment imposes excessive demands on the fitting capacity of models, consequently increasing the risk of overfitting. Moreover, as shown in Fig. 1(a), these methods commonly focus on encoding pixel-level information of the entire face associated with AU, disregarding the fact that capturing subtle feature differences when AUs activate is sufficient. As a result, these learning methods place significant demands on the model complexity to encode excessive redundant information, ultimately limiting the generalization ability of the model.
(a) Pixel-level feature learning methods
(b) Discriminative contrastive learning methods
@@
Representation learning approaches rooted in contrastive learning have demonstrated substantial promise [10] for classification. This approach aims to maximize the consistency of positive samples that are semantically similar while minimizing the consistency of negative samples that are semantically dissimilar [11]. By optimizing the contrastive loss, the mutual information between variables is either explicitly or implicitly optimized [12]. However, the direct integration of contrastive learning into AU detection encounters two principal challenges. The first challenge is the class imbalance issue of each AU type [7]. Most AU classes manifest a substantial imbalance between majority and minority class samples, with inactivated AUs vastly outnumbering their activated counterparts. This imbalance issue influences the classification decision threshold, potentially resulting in the direct classification of the majority of AUs as the negative class. The second challenge pertains to the noisy and false label problem [13]. Due to the subtle feature differences when AUs activate leading to the inherent difficulty in expertly labeling AUs, existing AU datasets may exhibit a few noisy and false labels, significantly undermining the overall fitting ability of models.
@@
To address these challenges, as shown in Fig. 1(b), we propose a discriminative contrastive learning framework designed to encode differential information between AUs from each sample pair, replacing the traditional pixel-level feature learning methods within the field of AU detection. The resultant loss is termed AUNCE, resembling the commonly utilized InfoNCE. Moreover, to mitigate the class imbalance issue of each AU type, we apply a negative sample re-weighting strategy, making modifications to the gradient magnitude of minority and majority class samples during back-propagation by adding importance weight to the denominator of the resultant loss AUNCE. Additionally, in response to the issue of noisy and false labeling in AUs, we introduce a positive sample samplying strategy to augment the supervision signal through self-supervised signals. This is achieved by establishing sampling ratios for various types of positive sample pairs, ensuring that more robust positive samples are derived from data augmentation with the self-supervised signal. The primary contributions of our work are succinctly summarized as follows:
1.
@@
We propose a discriminative contrastive learning framework to learn differential feature representations for AU detection, and the resultant loss is named AUNCE.
2.
@@
2 Related Work
We provide a comprehensive overview of two areas closely connected to our research: discriminative contrastive learning and contrastive learning for AU detection.
2.1 Discriminative contrastive learning
@@
In recent years, discriminative contrastive learning, a subset of deep self-supervised learning methods [15], has witnessed substantial progress. This approach adheres to the learn-to-compare paradigm [16], wherein a distinction is made between observed data and noisy or erroneous data, preventing the model from engaging in the reconstruction of pixel-level information. For example, Oord et al. [12] optimized feature extraction networks by maximizing consistency between predicted and actual outcomes of sequence data, introducing the widely-utilized InfoNCE loss, which implicitly or explicitly serves as a lower bound for mutual information. Chen et al. [11] proposes a prominent framework named SimCLR, which introduces a base encoder network and a projection head trained to maximize agreement utilizing the contrastive loss InfoNCE. Additionally, MoCo, proposed by He et al. [17], employs a momentum network to maintain a queue of negative samples for efficient contrastive learning. Chen and He [18] mitigate undesired collapsing solutions, eschewing the use of negative sample pairs through a stop-gradient operation. Despite variations and errors in representation encoders (denoted as ) and similarity measures across different tasks [19], the underlying concept remains consistent: training the representation encoder by optimizing the contrastive loss, achieved through contrasting similar pairs () and dissimilar pairs ().
@@
Apart from that, deep contrastive learning not only facilitates the automated extraction of refined feature representations through the comparison of positive and negative sample pairs but also incorporates diverse contrastive losses for optimization. Theoretical analyses concerning the generalization bounds of contrastive learning for classification tasks have been provided by Arora et al. [20]. Khosla et al. [21] extended the contrastive learning concept to supervised settings with a supervised contrastive learning loss, aiming to enhance feature expression using labeled data. Additionally, Chen et al. [22] proposed a semi-supervised contrastive learning algorithm, initially pretraining all data through contrastive learning and subsequently transferring knowledge to new models via distillation learning with the aid of labeled data.
@@
There is a discernible shift in the construction of sample pairs for contrastive learning. In cases where data is labeled or partially labeled, positive samples can be constructed using images belonging to the same category, rather than being confined to identical images. This approach augments the diversity of positive sample pairs, consequently enhancing feature expression. Furthermore, contrastive learning serves as a precursor to supervised learning, enabling model pretraining to significantly enhance the performance of subsequent supervised learning models. However, there has been no prior work that directly applies the InfoNCE-based discriminative contrastive learning framework to AU detection task based on static images, primarily due to inherent challenges in this field: class imbalance problem of each type of AUs, noisy and erroneous annotations, and difficulty in distinguishing hard negative samples.
2.2 Contrastive learning for AU detection
@@
Contrastive learning was firstly designed as a pretext task for unsupervised or self-supervised learning and was introduced to the domain of AU detection with limited annotations [23, 24, 25]. For example, Niu et al. [23] introduced a semi-supervised co-training approach that utilized contrastive loss to enforce conditional independence between facial representations generated by two Deep Neural Networks (DNNs). Chang et al. [24] present a knowledge-driven representation learning framework aimed for AU detection, incorporating a contrastive learning module that leverages dissimilarities among facial regions to effectively train the feature encoder. Additionally, Liu et al. [25] introduced a semi-supervised contrastive learning strategy that concurrently combined region learning, contrastive learning, and pseudo-labeling.
@@
Subsequent research endeavors extended contrastive learning to supervised settings, aiming at enhancing feature representation for AU detection through the utilization of labeled data [26, 27, 28]. Wu et al. [26] introduced a contrastive feature learning method, enabling a Convolutional Neural Network (CNN) to acquire distinctive features that represent differences between neutral and AU-activated facial images. Chen et al. [28] introduced a supervised hierarchical contrastive learning method (SupHCL) aimed at enhancing AU detection by fostering knowledge consistency across distinct facial images and various AUs. In a similar vein, Suresh et al. [27] proposed a feature-based positive matching contrastive loss, which facilitates learning distances between positives based on the similarity of corresponding AU embeddings.
However, the aforementioned methods mostly employed a combination of weighted cross-entropy loss and triplet loss to constrain the training process, which involves encoding pixel-level information of the entire face associated with AU, leading to a substantial amount of redundant information. To address this issue, Li et al. [13] formulated a self-supervised contrastive learning method, which introduced the discriminative contrastive learning framework MoCo [17] into dynamic AU detection task based on extensive amounts of unlabeled facial videos, which is aim to extract differential features between video frames, inspired by some discriminative contrastive learning techniques in video representation learning [29].
In summary, contrastive learning has demonstrated promising results in the field of AU detection. However, how to capture differential information between AUs from each sample pair instead of pixel-level information of the entire face associated with AU in AU detection task based on static images is still a problem that needs to be solved.
3 Method
@@
In Section 3, we present a discriminative contrastive learning framework specifically designed for the task of AU detection, aiming to incentivize the encoder to learn differential information between AUs from each sample pair, rather than focusing on pixel-level information of the entire face associated with AU. Subsequently, we confront two additional challenges: 1. Tackling the issue of class imbalance pertaining to each AU type in order to enhance the accuracy of our framework. 2. Addressing the problem of noisy and false labels, aiming to bolster the robustness of our models against erroneously annotated data.
3.1 Discriminative contrastive learning framework
@@
For the task of AU detection, which falls under the category of multi-label binary classification, the prevalent approach for training models has been the utilization of cross-entropy loss. This loss measures the disparity between predicted labels and ground truth labels. However, there exists variability in the occurrence rates of different AUs, with some AUs exhibiting low occurrence rates while others display high occurrence rates. Consequently, in contrast to the conventional cross-entropy loss, we employ a weighted cross-entropy loss [7] denoted as
@@
However, the classification methods based on cross-entropy loss suffer from a common issue of being overly constrained, which can lead to overfitting and limited generalization. This constraint becomes particularly problematic in the context of AU detection, where different AUs share a common encoder. For each AU of every sample, the predicted output must satisfy the aforementioned condition, placing a significant burden on the fitting capability of the encoder. To address this challenge, we introduce a discriminative contrastive learning framework for AU detection. As shown in Fig. 2. It comprises a positive sample mining module that integrates self-supervised signals with supervised signals, aiming to address the issue of noisy labels resulting from subtle differences in AUs. Additionally, a negative sample weighting module is included to tackle class imbalance and hard negative sample mining challenges. These modules will be discussed in detail in Sections 3.2 and 3.3, respectively. The resultant loss is formulated as:
@@
The contrastive loss distinguishes itself from weighted cross-entropy loss in the following aspects: 1. The weighted cross-entropy loss is computed based on the sample and its corresponding label , requiring the encoding of pixel-level information in order to align with the label. In contrast, the contrastive loss does not involve explicit labels but is computed based on the feature difference values between each pair of samples. This incentivizes the encoder to capture the differential information between AUs from each sample pair, without the need to focus on pixel-level information extraction of individual samples, reducing the borden of encoder. 2. The weighted cross-entropy loss imposes constraints on the absolute values of the predicted probabilities. However, for classification tasks, this is often unnecessary and can lead to overfitting as well as demanding higher fitting capability from the encoder. The contrastive loss focuses on constraining the similarity difference between positive and negative sample pairs without explicitly specifying their exact values. This is because the minimal value of is 0, which is achieved when for all negative sample index .
3.2 Negative sample re-weighting strategy
@@
Above equation follows the Boltzmann distribution. As the value of increases, the entropy of the distribution decreases significantly [31], leading to a sharper distribution in the region of high similarity. This has two consequences: 1. It imposes larger penalties on hard negative samples that are in close proximity to the anchor data point, enabling the model to learn more precise decision boundaries from subtle feature differences when AUs activate. 2. It controls the relative contribution of negative samples to the parameter updates, providing a means to address class imbalance issues.
@@
The primary distinction between our weighting method and HCL [30] lies in the fact that Eq. (3) does not divide by the mean as a normalization factor, enabling a more effective solution to the class imbalance issue in the field of AU detection. Specifically, when anchor data point belongs to the majority class, its positive sample belongs to the majority class, while its negative sample belongs to the minority class. In this case, we aim to have smaller gradient magnitudes for the positive samples (majority class) to update fewer parameters, while the negative samples (minority class) should have larger gradient magnitudes to update more parameters. To achieve this, we begin to set , which increases the impact of parameter updates caused by negative samples (minority class). On the other hand, when belongs to minority class, we begin to set to reduce the influence of parameter updates caused by majority class samples. By adjusting the value of , we can control the relative importance of positive and negative samples during the parameter updating, ensuring that the model pays more attention to minority class samples when necessary and balances the gradient magnitudes accordingly. This helps to balance the representation of different AUs in training set and improves the model ability to precisely classify all AU samples. If the mean were used as a normalization factor, the precise control over the parameter updates for majority and minority classes through the setting of would become unfeasible. This is because changes in would concurrently alter the mean, rendering the subsequent increase or decrease in weights after division by the mean uncertain. Therefore, the omission of division by the mean in Eq. (3), motivated by considerations of class imbalance, distinguishes our work from HCL.
3.3 Positive sample samplying strategy